OpenStudy (anonymous):

Factor completely. 8s^3+1 =

OpenStudy (anonymous):

\[8s^3+1=(2s)^3+1^3=(2s+1)((2s)^2-(2s)+1)=(2s+1)(4s^2-2s+1)\]